OpenStudy (anonymous):

Determine how many zeros are at the end of the number 132! show work and explanation.

OpenStudy (anonymous):

do you know what the greatest integer function is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the question boils down to how many factors of 5 are in 132, since there are more factors of 5 than of 2

OpenStudy (anonymous):

every time a 5 is paired with a 2 you will get 10, so that will contribute to a zero at the end of the numbers therefore solving this problem is the same as answering "how many factors of 5 are in 132!?"

OpenStudy (anonymous):

i dont know how i would show my work though. do 132 divided by 5?

OpenStudy (anonymous):

that is the first step you get 26.4 ignore the .4, keep the 26 then divide by \(5^2=25\)

OpenStudy (anonymous):

you get \(5.28\) ignore the .28, keep the 5

OpenStudy (anonymous):

then divide by \(5^3=125\) you get 1 and some remainder, keep the 1

OpenStudy (anonymous):

could yu show all the work at once please?

OpenStudy (anonymous):

@robtobey will show it maybe there is a snappier way, but i think that is basically what you have to do then add up those numbers

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

32 30*3+2 Sorry, no sophistication here, just brute force. Refer to the attachement.

OpenStudy (anonymous):

but how would i explain there are 32 zeros

OpenStudy (anonymous):

Refer to the attachment, the value of 132! calculated by Mathematica, and count them.

OpenStudy (anonymous):

yea but i ahve to show work or explain why

OpenStudy (anonymous):

\(26+5+1=32\)

OpenStudy (anonymous):

hope it is clear how i got those numbers